A 0.20 ml CO2 bubble in a cake batter is at 27°C. In the oven it gets heated to 177°C. What is its new volume?


1
Expert's answer
2021-11-11T19:13:02-0500

Apply

V1/T1 = V2/T2

0.2/(27+273) = V2/(177+273)

V2 = 0.2/(27+273)×(177+273)

V2 = 0.3 mL